# Memory Gym: Towards Endless Tasks to Benchmark Memory Capabilities of Agents

### v5

@@ 5.6 Investigating Transformer-XL’s Surprisingly Low Effectiveness
Figure 8: Experiments on Endless Mortar Mayhem with varied TrXL configurations. The learning rate (LR) is adjusted to decay from $2.75$e-$4$ to $1.0$e-$4$ over 160 million steps, compared to the previous $1.0$e-$5$. We also incorporate observation reconstruction (Obs. Rec.) and observe the greatest improvement based on a ground truth estimation head (GT) as sanity check. Another test augments TrXL’s query with absolute positional encoding (QPos). The final experiment combines the optimized learning rate, augmented query, and ground truth estimation.
The unexpectedly low performance of TrXL prompts us to examine several hypotheses. Endless Mortar Mayhem, presenting the largest performance gap, serves as our primary environment for these detailed investigations.
@@ 5.6.1 Inadequate Networks Capacity
Our TrXL baseline comprises 2.8 million trainable parameters, while GRU consists of 4.05 million, prompting the question of whether TrXL’s model architecture lacks the necessary capacity. To investigate this, we conducted experiments varying the number of layers (2, 3, 4), the embedding dimension size (256, 384, 512), and the memory window length (256, 384, 512). None of these adjustments led to closing the performance gap to GRU. Even when the GRU cell is scaled down to a dimension of 384, totaling 2.7 million parameters, it still surpasses TrXL by a large margin.
@@ 5.6.2 Weak Learning Signal
Figure 8 depicts several additional experiments aimed at amplifying the learning signal. We made a naive adjustment to the learning rate schedule, utilized observation reconstruction, and introduced a ground truth estimation head. This ground truth estimation head is added to the actor-critic model architecture and predicts the target position to move to. Labels are provided by the environment, and the estimates are optimized using the mean-squared error loss. While ground truth information is typically unavailable and considered inappropriate for this benchmark, it serves as a useful sanity check. It helps determine if an additional learning signal is advantageous in this context. When training incorporates ground truth estimation, there is a noticeable improvement in the agent’s policy. Previously, the IQM of completed commands stood at 18; with ground truth estimation, it reaches 50. This outcome implies that the model’s capacity might not be the primary constraint. However, using the ground truth estimation head can lead to instability in training, with the agent’s performance fluctuating between 20 and 50 completed commands after about 300 million steps. During the initial 160 million training steps, the learning rate linearly decays from $2.75$e-$4$ to $1.0$e-$5$. By setting the final decayed learning rate to $1.0$e-$4$, the IQM performance hits 24 commands. Lower performances are observed when adding the observation reconstruction head to either just TrXL or TrXL with the optimized learning rate schedule. In summary, augmenting the learning signal — either by adjusting the learning rate or by incorporating ground truth estimation — leads to more performant policies.
@@ 5.6.3 Lack of Temporal Information in the Initial Query
The next hypothesis pertains to the initial query of the first TrXL layer. This query is constructed solely based on the features extracted from the observation encoders at the current time step, encompassing only minimal temporal information. In contrast, the query of the second layer draws from the aggregated outcome of the memory window, thus capturing more substantial temporal information. However, we believe that the initial query could be further enriched with additional information. In Figure 8, one experiment augments the query with absolute positional encoding, providing direct access to the current time step. Despite this enhancement, the agent’s performance only moderately improves, reaching an IQM of 19 completed commands. We also explored embedding the original query with the previous output of the TrXL block. However, this simple approach did not yield positive results.
@@ 5.6.4 Harmful Off-Policy Data in the Episodic Memory
Next, it can be discussed whether the combination of PPO and the cached hidden states of TrXL’s episodic memory pose a threat. PPO is an on-policy algorithm that consumes the training data for a few epochs. Once the second training epoch on the same batch commences, the batch is already considered off-policy. However, PPO’s loss function can mitigate this issue and hence allows for several epochs. But this mitigation is likely limited and we wonder whether this issue is enlarged by the cached hidden states. In our current training setup, the agent collects 512 steps of data. If an episode exceeds 512 steps and is still running, the older hidden states from that episode become outdated over multiple epochs. To address this, future research should investigate the impact of stale hidden states on training efficacy. One potential strategy could involve computing more fresh hidden states alongside utilizing cached ones. Alternatively, cached hidden states could be periodically recomputed, albeit at the expense of increased inference overhead. Drawing parallels, Kapturowski et al. (2019) explored similar challenges with their recurrent and off-policy algorithm R2D2.
@@ 5.6.5 Indistinguishable Absolute Positional Encoding
The original work of TrXL introduces a relative positional encoding (Dai et al., 2019) as a consequence of extended context of lengths. However, when they refer to the absolute positional encoding, they only apply it to the range of the current window. If a window has a length of 4, the positions within the window are always the same (i.e. $[0,1,2,3]$), even though the window moves across the entire sequence. In our case, the absolute positional encoding considers the entire sequence. If the window has progressed in time, the applied positional encoding relies on the current time step (i.e. $[t-3,t-2,t-1,t]$).
Nevertheless, the absolute positional encoding needs to be defined prior training. When scaling it up, the question arises whether the agent’s model is capable of differentiating time steps, because a longer length indicates a more fine-grained sample frequency over the underlying positional encoding. To test this, we trained a TrXL agent on the finite task of Mortar Mayhem and configured an absolute positional encoding of length 1024 and 2048. Both experiments, which were repeated twice, resulted in an agent that poorly completes only 1 to 2 commands. This demands for questioning the absolute positional encoding in endless tasks. But if the elements of the absolute position encoding of length 2048 are not distinguishable, how do those agents master more than 20 commands in Endless Mortar Mayhem?
@@
If this assumption holds right that the absolute positional encoding is the culprit in our TrXL baseline, leveraging learned or relative positional encodings instead should improve performance. As seen in Figure 9, no notable advancement is accomplished by both variants. We observed two outliers among the 5 runs that utilized relative positional encoding. One underperformed, completing only 13 commands, while the other excelled, finishing more than 60 commands. Since the interquartile mean represents only the central 50% of the data, these outliers are not apparent in the sample efficiency curve. Given this discussion, a more in-depth exploration of positional encoding in this context is advisable.
@@ 5.7 Recurrence is not Vulnerable to Spotlight Perturbations
However, the results presented in Figure 6(g) neglect this hypothesis. This prompted us to conduct an ablation study, beginning with the previous hyperparameters. In this study, only a single hyperparameter or environment option was replaced with the current one at a time. Table 1 presents the results for Searing Spotlights when full observability is ensured. Initially, observation reconstruction was presumed to be the solution. However, the crucial adjustment proved to be the decision to avoid normalizing advantage estimates during optimization. Only the GRU agent, which omits advantage normalization, and the memory-less agent, based entirely on previous hyperparameters, recorded noteworthy success rates of 100% and 96%, respectively.
@@
Several comprehensive studies have delved into such nuances of PPO’s implementation (Engstrom et al., 2020; Andrychowicz et al., 2021; Huang et al., 2022). Among them, only Andrychowicz et al. (2021) turned off advantage normalization and found negligible performance impact. However, their extensive study focused on control tasks with undiscounted returns in the thousands, contrasting Memory Gym’s finite environments where the undiscounted returns do not exceed 2. During hyperparameter optimization, turning off advantage normalization also enhanced performance in Mortar Mayhem and Mystery Path. Given the critical importance of utilizing raw advantages, there remains an opportunity to delve even deeper into PPO to comprehend and refine its essential aspects.
@@ 6 Conclusion
In this study, we advanced Memory Gym’s environments to novel endless tasks, tailored to benchmark memory-based Deep Reinforcement Learning algorithms. These tasks feature a mounting challenge similar to the cumulative memory game “I packed my bag”. As the agent’s policy refines, the task dynamically expands, serving as an automatic curriculum. This innovative framework enables a thorough evaluation of memory-based agents, emphasizing their overall effectiveness beyond mere interaction efficiency with the environment. Our experimental results reveal that agents trained on these endless tasks consistently show stronger capabilities than those trained on finite tasks, demonstrating that finite environments offer only a restricted perspective on an agent’s full memory capabilities.
To foster a more inclusive research landscape, we contributed an open-source PPO baseline powered by Transformer-XL (TrXL). This baseline employs an attention mechanism applied to an episodic memory with a sliding window. When benchmarked against Memory Gym, the TrXL baseline competes head-to-head with another prominent baseline rooted in Gated Recurrent Unit (GRU). Notably, in finite environments, TrXL demonstrates superior effectiveness in Mortar Mayhem and enhanced sample efficiency in Mystery Path. However, in the Searing Spotlights environment, GRU emerges as the more efficient architecture. Perhaps our most unexpected revelation is the comeback of GRU in the endless environments. GRU surpassed TrXL by large margins, while also being computationally more efficient. Further probing into the Endless Mortar Mayhem environment revealed only marginal improvements for TrXL upon enriching the learning signal.
@@ Appendix C Results of Minor Baselines on the Finite Environments
This set of experiments reaffirms that agents trained on Memory Gym require memory. To show this, we train several naive baselines on simplified versions of the environments. These baselines include PPO without memory, PPO with frame stacking (4 and 16 grayscale frames), and PPO with absolute positional encoding (Vaswani et al., 2017) provided as vector observation. Note that the hyperparameters for these baselines are not tuned.
In Mortar Mayhem Act Grid, where the commands are completely provided as vector observation, PPO without memory is ineffective, while the frame stacking agents achieve nearly either 5 or 8 commands with a slight upward trend, while the one with absolute positional encoding follows closely, completing 9.5 commands. The positional encoding baseline adds the episode’s current time step to the agent’s observation allowing for temporal awareness within the sequential execution of commands. Once the complete task is present (Figure 10(b)), only GRU and TrXL prove to be effective.
We obtain a similar impression when training on Mystery Path Grid where the origin and goal are not perceivable by the agent (Figure 10(d)). If the origin and the goal are visible, a horizon of 16 frames is sufficient to train an effective policy as shown by the frame stacking agent (Figure 10(c)). Stacking 4 frames or leveraging positional encoding leads to an IQM success rate of 40% with a slight trend upward. The memory-less agent’s success rate goes up to about 11%.
No success is proven by the minor baselines in Searing Spotlights as seen in Figure 10(e). Only in this experiment, the naive baselines also leverage observation reconstruction.
@@ Appendix D GTrXL and LSTM Results on the Finite Environments
Figure 11 presents extended results for the finite environments. The performance metrics for TrXL and GRU agents remain consistent with those depicted in Figure 6. However, this figure displays twice as many data points on the x-axis. We include results from agents utilizing Gated Transformer-XL (GTrXL) (Parisotto et al., 2020) and the Long Short-Term Memory (LSTM) (Hochreiter and Schmidhuber, 1997). These models are benchmarked without any hyperparameter tuning. For GTrXL, biases of both 0 and 2 are tested, based on the suggestion by Parisotto et al. (2020) that a larger bias might enhance learning. All additional agents in the Searing Spotlights environment employ the observation reconstruction loss (Obs. Rec.), while it is omitted in the Mystery Path environment.
For Mortar Mayhem (as seen in Figure 11(a)), the LSTM agent settles at 3 commands, whereas the GTrXL agents using observation reconstruction both achieve an IQM of 2 commands. In the Mystery Path results (Figure 11(b)), the LSTM agent achieves full success, albeit requiring more samples compared to the GRU and TrXL agents. The GTrXL agent with a bias of 0 reaches a 94% success rate, while the one with a bias of 2 achieves only 81%, challenging the earlier assertion by Parisotto et al. (2020). The confidence interval for these results indicates potential training instability. In the Searing Spotlights environment, the LSTM agent with observation reconstruction is slightly less efficient than its GRU counterpart. However, GTrXL agents, though successful, lag behind TrXL with observation reconstruction.
It is important to note that we avoid making conclusive statements based on these results. This caution stems from the fact that the LSTM and GTrXL baselines are not tuned, and the GTrXL implementation awaits further verification.
@@ Appendix F Simulation Speed of the Environments
Memory Gym’s environments are implemented using Python and PyGame [^3] https://www.pygame.org. The simulation speeds are shown in Table 7. At first glance, the endless variants appear slower than the finite ones. This discrepancy is attributed to the shorter episodes in endless variants, which rely on less effective constant action policies, thereby increasing the impact of environment reset costs. Compared to related memory benchmarks, our environments perform exceptionally well, with only Procgen (Cobbe et al., 2020) being faster (Pleines et al., 2023). The training costs in our experiments are primarily due to the inference time of the baseline models.

### v6

@@ 5.6 Investigating Transformer-XL’s Surprisingly Low Effectiveness
Figure 8: Experiments on Endless Mortar Mayhem with varied TrXL configurations. The learning rate (LR) is adjusted to decay from $2.75$e-$4$ to $1.0$e-$4$ over 160 million steps, compared to the previous range from $2.75$e-$4$ to $1.0$e-$5$. We also incorporate observation reconstruction (Obs. Rec.) and observe the greatest improvement based on a ground truth estimation head (GT) as sanity check. Another test augments TrXL’s query with absolute positional encoding (QPos). The final experiment combines the optimized learning rate, augmented query, and ground truth estimation.
The unexpectedly low performance of TrXL prompts us to examine several hypotheses. Endless Mortar Mayhem, presenting the largest performance gap, serves as our primary environment for these investigations.
@@ 5.6.1 Inadequate Networks Capacity
Our TrXL baseline comprises 2.8 million trainable parameters, while GRU consists of 4.05 million, prompting the question of whether TrXL’s model architecture lacks the necessary capacity. To investigate this, we conducted experiments varying the number of layers (2, 3, 4), the embedding dimension size (256, 384, 512), and the memory window length (256, 384, 512). None of these adjustments led to closing the performance gap to GRU. Even when the GRU cell is scaled down to a dimension of 384, totaling 2.7 million parameters, it still surpasses TrXL by a large margin (Figure 8). Note that TrXL’s maximum context length, defined by the number of layers and window length (Equation 7), reaches 766 timesteps with 3 layers and a window length of 256. Therefore, the context length is not the limiting factor in any endless environment.
@@ 5.6.2 Weak Learning Signal
Figure 8 depicts several additional experiments aimed at amplifying the learning signal. We made a naive adjustment to the learning rate schedule, utilized observation reconstruction, and introduced a ground truth estimation head. This ground truth estimation head is added to the actor-critic model architecture and predicts the next target position to move to. Labels are provided by the environment, and the estimates are optimized using the mean-squared error loss. While ground truth information is typically unavailable and considered inappropriate for this benchmark, it serves as a useful sanity check, helping to determine if an additional learning signal is advantageous in this context.
When training incorporates ground truth estimation, there is improvement in the agent’s policy. Previously, the mean number of completed commands stood at 16; with ground truth estimation, it reaches a mean of up to 36 commands. As a default and during the initial 160 million training steps, the learning rate linearly decays from $2.75\times 1.0$e-$4$ to $1.0$e-$5$. By setting the final decayed learning rate to $1.0$e-$4$, the mean performance hits 25 commands. The results of leveraging ground truth estimation or tuning the learning rate schedule imply that the model’s current capacity is capable of retaining more information.
@@ 5.6.3 Lack of Temporal Information in the Initial Query
The next hypothesis pertains to the initial query of the first TrXL layer. This query is constructed solely based on the features extracted from the observation encoders at the current time step, encompassing only minimal temporal information. In contrast, the query of the second layer draws from the aggregated outcome of the memory window, thus capturing more substantial temporal information. However, we believe that the initial query could be further enriched with additional information. In Figure 8, one experiment augments the query with absolute positional encoding, providing direct access to the current time step. Despite this enhancement, the agent’s performance only moderately improves, reaching a mean of 18 completed commands. We also explored embedding the original query with the previous output of the TrXL block. However, this simple approach did not yield positive results.
None of the aforementioned measures individually reach the performance level of the GRU agent. However, when combined, they achieve a mean performance of 47 executed commands at best.
@@ 5.6.4 Harmful Off-Policy Data in the Episodic Memory
Next, it can be discussed whether the combination of PPO and the cached hidden states of TrXL’s episodic memory pose a threat. PPO is an on-policy algorithm that consumes the training data for a few epochs. Once the second training epoch on the same batch commences, the batch is already considered off-policy. However, PPO’s loss function can mitigate this issue and hence allows for several epochs. But this mitigation is likely limited and we wonder whether this issue is enlarged by the cached hidden states. In our current training setup, the agent collects 512 steps of data. If an episode exceeds 512 steps and is still running, the older hidden states from that episode become outdated over multiple iterations of PPO. To address this, future research should investigate the impact of stale hidden states on training efficacy or how one can solve the issue of truncated episodes during optimization. One potential strategy could involve computing more fresh hidden states alongside utilizing cached ones. Alternatively, cached hidden states could be periodically recomputed, albeit at the expense of increased inference overhead. Drawing parallels, Kapturowski et al. (2019) explored similar challenges with their recurrent, off-policy algorithm R2D2.
@@ 5.6.5 Indistinguishable Absolute Positional Encoding
The original work of TrXL introduces a relative positional encoding (Dai et al., 2019) as a consequence of extended context lengths. However, when they refer to the absolute positional encoding, they only apply it to the range of the current window. If a window has a length of 4, the positions within the window are always the same (i.e. $[0,1,2,3]$), even though the window moves across the entire sequence. In our case, the absolute positional encoding considers the entire sequence. If the window has progressed in time, the applied positional encoding relies on the current time step (i.e. $[t-3,t-2,t-1,t]$).
Nevertheless, the absolute positional encoding needs to be defined prior training. When scaling it up, the question arises whether the agent’s model is capable of differentiating time steps, because a longer length indicates a more fine-grained sample frequency over the underlying positional encoding. To test this, we trained a TrXL agent on the finite task of Mortar Mayhem and configured an absolute positional encoding of length 1024 and 2048. Both experiments, which were repeated twice, resulted in an agent that poorly completes only 1 to 2 commands. This demands for questioning the absolute positional encoding in endless tasks. But if the elements of the absolute position encoding of length 2048 are not distinguishable, how do those agents master more than 10 commands in Endless Mortar Mayhem?
@@
If this assumption is correct—that the absolute positional encoding is the culprit in our TrXL baseline—leveraging learned or relative positional encodings instead should improve performance. As seen in Figure 9, no notable advancement is accomplished by both variants. We observed two outliers among the 5 runs that utilized relative positional encoding. One underperformed, completing only 16 commands on average, while the other excelled, finishing 56 commands on average. Given this discussion, a more in-depth exploration of positional encoding in this context is advisable.
@@ 5.7 Recurrence is not Vulnerable to Spotlight Perturbations
However, the results presented in Figure 6(c) neglect this hypothesis. This prompted us to conduct an ablation study, beginning with the previous hyperparameters. In this study, only a single hyperparameter or environment option was replaced with the current one at a time. Table 1 presents the results for Searing Spotlights when full observability is ensured. Initially, observation reconstruction was presumed to be the solution. However, the crucial adjustment proved to be the decision to avoid normalizing advantage estimates during optimization. Only the GRU agent, which omits advantage normalization, and the memory-less agent, based entirely on previous hyperparameters, recorded noteworthy success rates of 100% and 96%, respectively.
@@
Several comprehensive studies have delved into such nuances of PPO’s implementation (Engstrom et al., 2020; Andrychowicz et al., 2021; Huang et al., 2022a). Among them, only Andrychowicz et al. (2021) turned off advantage normalization and found negligible performance impact. However, their extensive study focused on control tasks with undiscounted returns in the thousands, contrasting Memory Gym’s finite environments where the undiscounted returns do not exceed 2. During hyperparameter optimization, turning off advantage normalization also enhanced performance in Mortar Mayhem and Mystery Path. Given the critical importance of utilizing raw advantages, there remains an opportunity to delve even deeper into PPO to comprehend and refine its essential aspects.
@@ 6 Conclusion
In this study, we advanced Memory Gym’s environments to novel endless tasks, tailored to benchmark memory-based Deep Reinforcement Learning algorithms. These tasks feature a mounting challenge similar to the cumulative memory game “I packed my bag”. As the agent’s policy refines, the task dynamically expands, serving as an automatic curriculum. This innovative framework enables a thorough evaluation of memory-based agents, emphasizing their overall effectiveness beyond mere interaction efficiency with the environment. Our experimental results reveal that agents trained on these endless tasks consistently show stronger capabilities than those trained on finite tasks, demonstrating that finite environments offer only a truncated view on an agent’s full memory capabilities.
To foster a more inclusive research landscape, we contributed an open-source PPO baseline powered by Transformer-XL (TrXL). This baseline employs an attention mechanism applied to an episodic memory with a sliding window. When benchmarked against Memory Gym, the TrXL baseline competes head-to-head with another prominent baseline found in Gated Recurrent Unit (GRU). Notably, in finite environments, TrXL is more effective than GRU, but strongly depends on the observation reconstruction loss. Our most unexpected revelation is the comeback of GRU in the endless environments. GRU surpasses TrXL by large margins, while also being computationally more efficient. Further probing into the Endless Mortar Mayhem environment revealed only marginal improvements for TrXL upon enriching the learning signal.
@@ Appendix C Results of Minor Baselines on the Finite Environments
This set of experiments reaffirms that agents trained on Memory Gym require memory. To show this, we train several naive baselines on simplified versions of the environments. These baselines include PPO without memory, PPO with frame stacking (4 and 16 grayscale frames), and PPO with absolute positional encoding (Vaswani et al., 2017) provided as vector observation. Note that the hyperparameters for these baselines are not tuned. Figure 10 shows the mean and standard deviation of 5 independent seeds for each presented agent.
In Mortar Mayhem Act Grid, where the commands are completely provided as vector observation, PPO without memory is ineffective, while the frame stacking agents achieve nearly either 4.7 or 7.3 commands with a slight upward trend, while the one with absolute positional encoding follows closely, completing 9.6 commands. The positional encoding baseline adds the episode’s current time step to the agent’s observation allowing for temporal awareness within the sequential execution of commands. Once the complete task is present (Figure 10(b)), only GRU and TrXL prove to be effective.
We obtain a similar impression when training on Mystery Path Grid where the origin and goal are not perceivable by the agent (Figure 10(d)). If the origin and the goal are visible, a horizon of 16 frames is sufficient to train an effective policy as shown by the frame stacking agent (Figure 10(c)). Stacking 4 frames or leveraging positional encoding leads to a success rate of 42% and 49% respectively. The memory-less agent’s rate goes up to about 29%.
Limited success is proven by the minor baselines in Searing Spotlights as seen in Figure 10(e). Only in this experiment, the naive baselines also leverage observation reconstruction.
@@ Appendix D GTrXL and LSTM Results on the Finite Environments
Figure 11 presents extended results for the finite environments, showing the mean and standard deviation across 5 independent runs for each trained agent. The performance metrics for TrXL and GRU agents remain consistent with those depicted in Figure 6. We include results from agents utilizing Gated Transformer-XL (GTrXL) (Parisotto et al., 2020) and the Long Short-Term Memory (LSTM) (Hochreiter and Schmidhuber, 1997). These models are benchmarked without any hyperparameter tuning. For GTrXL, biases of both 0 and 2 are tested, based on the suggestion by Parisotto et al. (2020) that a larger bias might enhance learning. All additional agents in the Searing Spotlights environment employ the observation reconstruction loss (Obs. Rec.), while it is omitted in the Mystery Path environment.
For Mortar Mayhem (as seen in Figure 11(a)), the LSTM agent settles at an average of 2.2 commands, whereas the GTrXL agents achieve a mean of 2.1 and 2.8 commands. In the Mystery Path results (Figure 11(b)), the LSTM agent completes 90% of the task. The GTrXL agent with a bias of 0 reaches a 88% progression, while the one with a bias of 2 achieves 86%, challenging the earlier assertion by Parisotto et al. (2020). In the Searing Spotlights environment, the LSTM agent with observation reconstruction is slightly less efficient than its GRU counterpart. However, GTrXL agents, though still growing to success, lag behind TrXL with observation reconstruction.
We avoid making conclusive statements based on these results. This caution stems from the fact that the LSTM and GTrXL baselines are not tuned.
@@ Appendix F Simulation Speed of the Environments
Memory Gym’s environments are implemented using Python and PyGame [^3] https://www.pygame.org. The simulation speeds are shown in Table 7. At first glance, the endless variants appear slower than the finite ones. This discrepancy is attributed to the shorter episodes in endless variants, which rely on less effective constant action policies, thereby increasing the impact of environment reset costs. Compared to related memory benchmarks, our environments perform well, with only Procgen (Cobbe et al., 2020) being faster (Pleines et al., 2023). The training costs in our experiments are primarily due to the inference time of the baseline models.